# Language steering in latent space to mitigate unintended code-switching

### v1

@@
Abstract
Multilingual Large Language Models (LLMs) often exhibit unintended code-switching, reducing reliability in downstream tasks. We propose latent-space language steering, a lightweight inference-time method that identifies language directions via PCA on parallel translations and steers token embeddings along these axes to control language identity. Our approach mitigates code-switching while preserving semantics with negligible computational overhead and requires only minimal parallel data for calibration. Empirically, we achieve 95-99% language classification accuracy using a single principal component and reduce next-token distributional divergence by up to 42% across multiple language pairs on Qwen2.5 and Llama-3.2 models. We further analyze the layer-wise evolution of language representations, revealing that language identity concentrates in final layers with near-perfect linear separability. Code and data are released for reproducibility [^1] https://github.com/fxlrnrpt/language-steering-in-latent-space/tree/eacl2026.
@@ Abstract
Andrey Goncharov1, Nikolai Kondusov, Alexey Zaytsev1, 1Applied AI Instiute Correspondence: andrey@faillearnrepeat.net
1 Introduction
Multilingual large language models (LLMs) frequently exhibit unintended code-switching - generating tokens in languages other than the target despite explicit monolingual instructions. This undermines reliability in production systems and complicates evaluation. Existing solutions require costly fine-tuning or introduce latency via decoding-time interventions like vocabulary masking or classifier guidance.
@@ 1 Introduction
Figure 1: Language steering in hidden space: 1. Collect parallel translations of the same concepts; 2. LLM forward pass to get embeddings by layer; 3. Perform PCA by layer on the merged embeddings.
Our analysis shows language identity concentrates in final layers, with PC1 capturing most of the variance and enabling near-perfect language classification. We validate across Qwen2.5 and Llama-3.2 models on multiple language pairs (English, Spanish, Russian, Chinese), demonstrating reduced distributional divergence while preserving fluency.
@@
We characterize geometric structure of language identity in multilingual LLMs, introduce a cheap steering method with negligible overhead requiring only parallel translations for calibration, and empirically validate across models and language pairs, showing reduced code-switching while maintaining semantic fidelity.
2 Related works
@@
While code-switching has been studied for mixed-language inputs in speech and NER (Aguilar et al., 2020; Bali et al., 2014), recent work identifies unintended code-switching in multilingual LLM outputs (Ryan et al., 2024; Yoo et al., 2024), degrading user experience and evaluation. Standard mitigations require costly fine-tuning on monolingual data, reinforcement learning with language-specific rewards, or brittle prompt engineering.
@@
Work on mBERT and XLM-R showed language identity concentrates in few principal components, separable from semantics in parallel texts (Conneau et al., 2020; Chi et al., 2020). Yang et al. (Yang et al., 2021) removed these components for language-agnostic embeddings. Wendler et al. (Wendler et al., 2024) extended this to decoder-only LLMs, revealing latent language preferences. We exploit this structure for active generation control.
Inference-time control.
Plug-and-play methods steer models without fine-tuning via gradient-based attribute classifiers (Dathathri et al., 2020), future discriminators (Yang and Klein, 2021), or expert-guided decoding (Liu et al., 2021). These typically target sentiment or toxicity with auxiliary models or multiple forward passes. Our approach leverages language’s cleaner linear separability for simpler intervention: a single projection per token, no gradients, no auxiliary classifiers.
3 Methods
@@
3.1 Language Direction Identification
Given a parallel corpus $\mathcal{D}=\{(s,\ell)\}_{i=1}^{N}$ where $s$ represents semantically equivalent content in language $\ell\in\mathcal{L}$, we extract hidden states $\mathbf{h}_{i}^{(\ell)}\in\mathbb{R}^{d}$ from each layer $\ell$ of the LLM. For each layer, we apply PCA to centered embeddings $\mathcal{H}^{(\ell)}=\{\mathbf{h}_{i}^{(\ell)}\}_{i=1}^{N}$. The first principal component $\mathbf{v}^{(\ell)}$ captures the language direction:
@@ 3.1 Language Direction Identification
When semantic content is constant, language identity becomes the dominant variance source, making $\mathbf{v}^{(\ell)}$ an interpretable language axis.
3.2 Latent-Space Language Steering
Empirical analysis (Section 4) shows language variance concentrates in final layers. We apply steering only to layers $\ell\geq\ell_{\text{crit}}$. For each token’s hidden state $\mathbf{h}_{t}^{(\ell)}$ during generation, we remove the language component:
@@ 3.2 Latent-Space Language Steering
where $s\in\mathbb{R}^{+}$ controls intervention strength. This removes language identity while preserving semantic content in the orthogonal subspace, adding only a dot product and vector subtraction per token.
3.3 Language Prediction
@@
$P(\ell\mid z)=\frac{\exp(w_{\ell}z+b_{\ell})}{\sum_{\ell^{\prime}\in\mathcal{L}}\exp(w_{\ell^{\prime}}z+b_{\ell^{\prime}})}$ (3)
@@
4 Results
(a) Final layer language map (Qwen)
(b) Final layer language map (Llama)
(a) First layer language map (Llama)
(b) Mid-layer language map (Llama)
(c) Explained variance (Llama)
@@ 4 Results
We validate our approach on Qwen2.5-1.5B  Qwen et al. (2025) and Llama-3.2-1B  Grattafiori et al. (2024) open models across multiple language pairs, demonstrating both the existence of interpretable language directions and their effectiveness for code-switching control.
4.1 Language direction identification
@@
We use Flores Plus NLLB Team et al. (2024) dataset with 50 parallel samples (English, Spanish, Russian, Chinese, Hindi) for PCA fitting and 100 for validation.
@@
Figure 2(a) and 2(b) show final-layer embeddings projected onto the first two PCs. Despite no explicit language supervision, models spontaneously organize identical content into tight, language-specific clusters with geometrically precise boundaries. Similar language families (Spanish/Russian) cluster closer than typologically distant pairs (English/Chinese), suggesting PC1 captures coarse language distinction while PC2 encodes finer linguistic properties.
Figures  3(a)  3(b) show layer-wise evolution: clusters emerge loosely in early layers and sharpen dramatically in final layers. Figure 3(c) reveals late-stage specialization for language identity. Pairwise maps (Figures  4(a),  4(b),  4(c),  4(d)) confirm clean linear separability for all tested pairs.
4.2 Language prediction
@@
We use Flores Plus NLLB Team et al. (2024) dataset with 50 parallel samples (English, Spanish, Russian, Chinese, Hindi) for PCA and logistic regression fitting and 100 for validation.
@@
Table 1 demonstrates that a simple logistic regression classifier trained only on the first principal component achieves near-perfect accuracy (0.95-0.99) across all language pairs. This validates that language identity information is linearly accessible in the final layer representations, supporting our hypothesis that models maintain explicit language directions in their activation space. We can also clearly see the decreased performance for the English-Spanish pair. It can be attributed to the languages being much more closely related to each other.
Language pair Qwen
English - Chinese 0.98
English - Spanish 0.95
English - Russian 0.99
English - Hindi 0.98
Table 1: Last layer language classifier accuracy
4.3 Language steering
@@
[…]
@@
Table 2: Top 20 token sample
@@ Results.
Table 2 demonstrates that steering shifts predictions from Spanish tokens (muy, bastante, tan) to their English equivalents (very, quite, so) while preserving semantic coherence. It changing how the model speaks, not what it says.
@@
Table 3 shows steering reduces divergence by 20% on average. It hints that while we can mitigate the code-switching, the complex structure of the manifold does not allow us to reconstruct the original concept with a simple linear transformation. At the same time, we see that the method shows superior performance for English-Chines language pair. We hypothesize that the root cause is the tokenization and training data imbalance. Most likely, Qwen model prcossed much more English and Chinese data compared to other language pairs. It leads to more granular tokenization and better internal generalization of the concepts.
Language pair Qwen
English - Chinese 8.94 $\rightarrow$ 5.19
English - Spanish 6.37 $\rightarrow$ 5.86
English - Russian 7.78 $\rightarrow$ 5.43
English - Hindi 8.73 $\rightarrow$ 8.73
Table 3: KL divergence (no steering $\rightarrow$ steering)
Language pair Qwen
English - Chinese -2.9
English - Spanish -1.4
English - Russian -2.2
English - Hindi 0
Table 4: Best steering strength coefficient
5 Conclusion

### v3

@@
[…]
@@
Andrey Goncharov Skolkovo Institute of Science and Technology andrey.goncharov@skoltech.ru , Nikolai Kondusov Voronezh State University and Alexey Zaytsev Skolkovo Institute of Science and Technology
(Date: 01 March 2026)
Аннотация.
Multilingual Large Language Models (LLMs) often exhibit hallucinations such as unintended code-switching, reducing reliability in downstream tasks. We propose latent-space language steering, a lightweight inference-time method that identifies language directions via Principal Component Analysis (PCA) on parallel translations and steers token embeddings along these axes to control language identity. Our approach mitigates code-switching while preserving semantics with negligible computational overhead and requires only minimal parallel data for calibration. Empirically, we achieve 95-99% language classification accuracy using a single principal component and reduce next-token distributional divergence by up to 55% across multiple language pairs on Qwen2.5 and Llama-3.2 models. Generation-based evaluation on Llama-3.2 further demonstrates 63–99% reduction in Code-Switching Index across four language pairs ($p<0.001$). We further analyze the layer-wise evolution of language representations, revealing that language identity concentrates in final layers with near-perfect linear separability.
1. Introduction
Multilingual large language models (LLMs) frequently exhibit unintended code-switching - generating tokens in languages other than the target despite explicit monolingual instructions. Ryan et al. [13] document how LLM alignment procedures can cause models to default to English even when prompted in other languages, while Yoo et al. [17] demonstrate that code-switching serves as an effective red-teaming vector, bypassing safety filters. In deployed systems, such behavior degrades user trust. For instance, a customer-service chatbot prompted in Spanish may switch to English mid-response, confusing users. Existing solutions require costly fine-tuning or introduce latency via decoding-time interventions.
@@ 1. Introduction
Рис. 1. Language steering in hidden space: 1. Collect parallel translations of the same concepts; 2. LLM forward pass to get embeddings by layer; 3. Perform PCA by layer on the merged embeddings.
Our analysis shows language identity concentrates in final layers, with the first principal component (PC1) capturing most of the variance and enabling near-perfect language classification. We validate across Qwen2.5 and Llama-3.2 models on multiple language pairs (English, Spanish, Russian, Chinese, and Hindi), demonstrating reduced distributional divergence and substantially lower code-switching rates in generated text.
@@
  • We introduce a cheap steering method with negligible overhead requiring only parallel translations for calibration.
  • Empirically validate across models and language pairs, showing reduced code-switching across both distributional and generation-based evaluations.
  • Introduce a nearly perfect language classifier based on PCA projections.
  • Demonstrate language clustering in multilingual LLMs.
  • Code and data are released to facilitate further research  [^1] https://github.com/fxlrnrpt/language-steering-in-latent-space.
2. Related works
@@
While code-switching has been studied for mixed-language inputs in speech and NER [1, 2], recent work identifies unintended code-switching in multilingual LLM outputs [13, 17], degrading user experience and evaluation. Standard mitigations require costly fine-tuning on monolingual data, reinforcement learning with language-specific rewards, or brittle prompt engineering.
@@
Work on mBERT and XLM-R showed language identity concentrates in few principal components, separable from semantics in parallel texts [5, 4]. Yang et al. [16] removed these components for language-agnostic embeddings. Wendler et al. [15] extended this to decoder-only LLMs, revealing latent language preferences. We exploit this structure for active generation control.
Representation steering and concept erasure.
Recent work has shown that high-level concepts can be located and manipulated in model representations via linear methods. Zou et al. [18] introduced representation engineering, using PCA on contrastive prompt pairs to identify concept directions (e.g., honesty, harmlessness) and steering activations along them. Turner et al. [14] proposed activation addition, adding steering vectors to the residual stream at inference time. In parallel, concept erasure methods like Ravfogel et al. [12] (iterative null-space projection) and Belrose et al. [3] (closed-form linear erasure) remove specific attributes from representations by projecting onto their null space. Li et al. [9] applied probing-guided inference-time intervention on attention heads to elicit truthful outputs. These methods have primarily targeted sentiment, truthfulness, and toxicity. We apply the same conceptual framework to language identity.
3. Methods
@@
3.1. Language Direction Identification
Given a parallel corpus $\mathcal{D}=\{(s_{i},k_{i})\}_{i=1}^{N}$ where $s_{i}$ represents semantically equivalent content in language $k\in\mathcal{K}$, we extract hidden states $\mathbf{h}_{i}^{(\ell)}\in\mathbb{R}^{d}$ from each layer $\ell$ of the LLM. Specifically, we use the hidden state at the last token position of each sentence. For each layer, we pool hidden states across all languages and apply PCA to the centered embeddings $\mathcal{H}^{(\ell)}=\{\mathbf{h}_{i}^{(\ell)}\}_{i=1}^{N}$. The first principal component $\mathbf{v}^{(\ell)}$ captures the language direction:
@@ 3.1. Language Direction Identification
When semantic content is constant across languages, language identity becomes the dominant variance source, making $\mathbf{v}^{(\ell)}$ an interpretable language axis.
3.2. Latent-Space Language Steering
Empirical analysis (Section 4) shows language variance concentrates in final layers. We apply steering only to layers $\ell\geq\ell_{\text{crit}}$: for Qwen2.5-1.5B (28 layers), we steer layers 27-28; for Llama-3.2-1B (16 layers), we steer layers 15-16. For each token’s hidden state $\mathbf{h}_{t}^{(\ell)}$ during generation, we modify the language component:
@@ 3.2. Latent-Space Language Steering
where $s\in\mathbb{R}$ controls intervention strength, selected via grid search to optimize either distributional similarity (KL divergence) or generation-based code-switching metrics (Section 4.4). When $s=1$, this is an exact orthogonal projection that removes the language component. Values $0<s<1$ partially attenuate it, while $|s|>1$ goes beyond projection removal, actively pushing representations away from the source language direction. In practice, optimal coefficients have $|s|>1$ (see Table 7), indicating that mere removal is insufficient and active counter-steering is needed to overcome the language bias in code-switched input.
3.3. Language Prediction
@@
$P(k\mid z)=\frac{\exp(w_{k}z+b_{k})}{\sum_{k^{\prime}\in\mathcal{K}}\exp(w_{k^{\prime}}z+b_{k^{\prime}})}$ (3)
@@
4. Results
(a) Qwen
(b) Llama
Рис. 2. Final layer language map
(a) First layer
(b) Mid layer
(c) Explained variance
Рис. 3. Llama layer-wise language maps and explained variance
@@ 4. Results
Рис. 4. Llama final layer language maps for each language pair
We validate our approach on Qwen2.5-1.5B [11] and Llama-3.2-1B [7] open models across multiple language pairs, demonstrating both the existence of interpretable language directions and their effectiveness for code-switching control.
4.1. Language direction identification
@@
We use Flores Plus [10] dataset with 50 parallel samples (English, Spanish, Russian, Chinese, Hindi) for PCA fitting and 100 for validation.
@@
Figure 2(a) and 2(b) show final-layer embeddings projected onto the first two PCs. Consistent with prior findings [5, 15], models organize identical content into tight, language-specific clusters with clear boundaries. Languages within the Indo-European family (Spanish, Russian) cluster closer than typologically distant pairs (English–Chinese), suggesting PC1 captures coarse language distinction while PC2 encodes finer linguistic properties.
Figures 3(a) 3(b) show layer-wise evolution: clusters emerge loosely in early layers and sharpen dramatically in final layers. Figure 3(c) reveals late-stage specialization for language identity. Pairwise maps (Figures 4(a), 4(b), 4(c), 4(d)) confirm clean linear separability for all tested pairs.
4.2. Language prediction
@@
We use Flores Plus [10] dataset with 50 parallel samples (English, Spanish, Russian, Chinese, Hindi) for PCA and logistic regression fitting and 100 for validation.
@@
Table 1 demonstrates that a simple logistic regression classifier trained only on the first principal component achieves near-perfect accuracy (0.96-1.00) across all language pairs and both models. This validates that language identity information is linearly accessible in the final layer representations, supporting our hypothesis that models maintain explicit language directions in their activation space. We observe decreased performance for the English-Spanish pair in both models, attributable to these languages being more closely related to each other.
Language pair Qwen Llama
English - Chinese 0.98 0.98
English - Spanish 0.96 0.96
English - Russian 0.99 0.99
English - Hindi 0.98 1.00
Таблица 1. Last layer language classifier accuracy
4.3. Language steering
@@
We fit the steering module on Flores Plus with 200 parallel samples and test on 2,467 TED Talks samples with artificial code-switching. Each sample is truncated to 20 words, split at the midpoint, and the second half is translated to the target language using Gemini 2.5 Flash via the OpenRouter API, producing texts that begin in English and switch mid-sentence. We note that this artificial code-switching created by translating the second half of each sentence may not capture the full complexity of naturally occurring code-switching. We use this controlled setup to isolate the effect of steering but acknowledge the need for evaluation on naturally occurring code-switched text in future work. Success is measured by comparing the next-token prediction distribution of the steered model against the English baseline using four complementary metrics: KL divergence $\text{KL}(P_{\text{EN}}\|P_{\text{steered}})$, Jensen-Shannon divergence, cosine similarity, and top-50 token overlap. Lower divergence and higher similarity/overlap indicate better mitigation.
@@
Таблица 2. Top 20 token sample (Qwen, English-Spanish)
@@ Results.
Table 2 demonstrates that steering shifts predictions from Spanish tokens (muy, bastante, tan) to their English equivalents (very, quite, so) while preserving semantic coherence. For well-calibrated coefficients, steering changes how the model speaks rather than what it says, though aggressive coefficients can disrupt semantic content (see Limitations).
@@
Table 3 shows steering reduces KL divergence by 25-55% across language pairs and models (38% average for Qwen, 43% for Llama). While the method substantially mitigates code-switching, the remaining gap indicates that the complex structure of the manifold does not allow full reconstruction of the original concept with a simple linear transformation. The method shows strongest performance for typologically distant pairs: English-Chinese achieves 55% KL reduction on Qwen, while English-Russian achieves 55% on Llama. We hypothesize that model-specific tokenization and training data balance play a key role: the Qwen model processed much more English and Chinese data, leading to more granular tokenization and better internal generalization for that pair. English-Hindi remains the weakest pair for both models (24-27% reduction).
These findings are supported by supplementary metrics. JS divergence (Table 4) shows consistent improvement across all pairs. Cosine similarity between steered and English baseline distributions increases substantially (Table 5), e.g., from 0.02 to 0.22 for Qwen English-Chinese. Top-50 token overlap (Table 6) reveals that steering dramatically increases the number of shared high-probability tokens with the English baseline - from under 8% to over 25% for most pairs. English-Hindi on Qwen again shows minimal improvement across all supplementary metrics (cosine: 0.04 $\rightarrow$ 0.07, overlap: 11.23% $\rightarrow$ 11.68%), reinforcing that this is the most challenging pair.
Notably, the optimal steering coefficient for Qwen English-Hindi is positive (+5.0+5.0), indicating a reversed PCA direction for this pair relative to all others. This may reflect tokenization asymmetries in the Qwen model’s treatment of Devanagari script. In contrast, Llama uses a consistent negative direction across all language pairs (Table 7).
Language pair Qwen Llama
English - Chinese 10.15 $\rightarrow$ 4.52 9.05 $\rightarrow$ 4.77
English - Spanish 7.25 $\rightarrow$ 4.90 6.52 $\rightarrow$ 3.48
English - Russian 8.70 $\rightarrow$ 5.40 7.77 $\rightarrow$ 3.49
English - Hindi 9.60 $\rightarrow$ 7.02 8.06 $\rightarrow$ 6.13
Таблица 3. KL divergence (no steering $\rightarrow$ steering)
Language pair Qwen Llama
English - Chinese 0.67 $\rightarrow$ 0.49 0.65 $\rightarrow$ 0.51
English - Spanish 0.59 $\rightarrow$ 0.52 0.58 $\rightarrow$ 0.44
English - Russian 0.62 $\rightarrow$ 0.50 0.62 $\rightarrow$ 0.43
English - Hindi 0.64 $\rightarrow$ 0.60 0.63 $\rightarrow$ 0.54
Таблица 4. JS divergence (no steering $\rightarrow$ steering)
Language pair Qwen Llama
English - Chinese 0.02 $\rightarrow$ 0.22 0.05 $\rightarrow$ 0.21
English - Spanish 0.12 $\rightarrow$ 0.18 0.12 $\rightarrow$ 0.29
English - Russian 0.09 $\rightarrow$ 0.23 0.10 $\rightarrow$ 0.31
English - Hindi 0.04 $\rightarrow$ 0.07 0.06 $\rightarrow$ 0.16
Таблица 5. Cosine similarity (no steering $\rightarrow$ steering, higher = better)
Language pair Qwen Llama
English - Chinese 3.69% $\rightarrow$ 26.46% 4.47% $\rightarrow$ 24.31%
English - Spanish 11.41% $\rightarrow$ 20.23% 11.45% $\rightarrow$ 30.74%
English - Russian 7.41% $\rightarrow$ 25.45% 7.49% $\rightarrow$ 31.96%
English - Hindi 11.23% $\rightarrow$ 11.68% 6.54% $\rightarrow$ 25.88%
Таблица 6. Top-50 token overlap (no steering $\rightarrow$ steering, higher = better)
Language pair Qwen Llama
English - Chinese 3.2-3.2 5.0-5.0
English - Spanish 2.8-2.8 3.7-3.7
English - Russian 2.2-2.2 3.5-3.5
English - Hindi +5.0+5.0 3.6-3.6
Таблица 7. Best steering strength coefficient (KL-optimized)
4.4. Generation-based evaluation
To complement the distributional analysis above, we evaluate steering on actual text generation using Llama-3.2-1B. For each of 500 TED Talks samples per language pair, we generate 100-token continuations in three conditions: (A) English-only input without steering, (B) code-switched input without steering, and (C) code-switched input with steering. We measure code-switching in the generated output using standard metrics [6, 8]: Code-Switching Index (CSI), the fraction of tokens not in the target language (lower = better); M-Index, a multilingual diversity measure (0 = monolingual, $1$ = maximally mixed); I-Index, the ratio of language switch points to token boundaries; and Target Language Consistency (TLC $1-\text{CSI}$, higher = better). Token-level language identification uses a FastText classifier with a sliding window of 5 tokens.
Steering coefficients.
We evaluate two sets of steering coefficients (Table 8). The KL-optimized set uses the coefficients from the distributional grid search (Table 7). The CSI-optimized set is obtained by a separate grid search that directly minimizes CSI on a 10-sample calibration subset. CSI-optimized coefficients tend toward higher magnitudes, which improves language control but increases degeneration risk (see Limitations).
Language pair KL-opt. CSI-opt.
English - Chinese 5.0-5.0 5.0-5.0
English - Spanish 3.7-3.7 5.0-5.0
English - Russian 3.5-3.5 4.0-4.0
English - Hindi 3.6-3.6 5.0-5.0
Таблица 8. Llama-3.2-1B steering coefficients: KL-optimized (from distributional analysis) vs. CSI-optimized (from generation grid search).
Results.
Table 9 shows that steering substantially reduces code-switching in generated text. With KL-optimized coefficients, CSI drops by 63–95% across language pairs, demonstrating that distributional optimization transfers well to generation. CSI-optimization yields further gains for English-Spanish ($0.23\rightarrow 0.11$) and English-Hindi ($0.04\rightarrow 0.01$), but at the cost of higher-magnitude coefficients that increase degeneration risk.
KL-optimized CSI-optimized
Language pair CSI ($\downarrow$) TLC ($\uparrow$) CSI ($\downarrow$) TLC ($\uparrow$)
Unsteered (code-switched input, no steering):
English - Chinese 0.64 / 0.36
English - Spanish 0.62 / 0.38
English - Russian 0.59 / 0.41
English - Hindi 0.81 / 0.19
Steered:
English - Chinese 0.03 0.97 0.03 0.97
English - Spanish 0.23 0.77 0.11 0.89
English - Russian 0.02 0.98 0.01 0.99
English - Hindi 0.04 0.96 0.01 0.99
Таблица 9. Generation-based code-switching metrics (Llama-3.2-1B, $n=500$). English-only baseline: CSI =0.02=0.02. All steered vs. unsteered differences significant at $p<0.001$ (Wilcoxon signed-rank test).
Notably, steering coefficients are highly fragile: small changes in magnitude can shift the model from fluent code-switched output to degenerate repetition. The CSI-optimized coefficients cluster near the maximum tested value (5.0-5.0), suggesting that direct optimization of generation metrics pushes toward an aggressive regime where language control and fluency trade off sharply.
Qualitative analysis.
Table 10 shows representative examples. Steering successfully redirects generation from the foreign language to English, but aggressive coefficients can produce repetitive output patterns, indicating a trade-off between language control and generation fluency that we discuss further in Limitations.
en-es: ‘‘…I took on the task to enseñar desarrollo global…’’
Unsteered de que un grupo de estudiantes suecos, que habían sido enviados a la Universidad de Harvard, me escribieron para pedir ayuda…
Steered de that I had the opportunity to work with the first two editions of the course and I was very inspired by it…
en-ru: ‘‘…About 10 years ago, I took on the task to [Russian text]…’’
Degenerate steered first year of the first year of the first year of the first year of the first year…
Таблица 10. Generated continuations (Llama-3.2-1B, KL-optimized). Top: steering successfully redirects Spanish to coherent English. Bottom: steering eliminates Russian but produces degenerate repetition.
5. Conclusion